# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ 4.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
In the second set of experiments, we empirically validate the results of Theorem 3.3 and Lemma 3.4 on the ModelNet40 and Polarity datasets. We build neural networks that consist of three layers: (i) a fully-connected layer; (ii) an aggregation function; and (iii) a second fully-connected layer. Therefore, those models first transform the elements of the input multisets using an affine function, then aggregate the representations of the elements of each multiset and finally they transform the aggregated representations using a another affine function. Note that the Lipschitz constant of an affine function is equal to the largest singular value of the associated weight matrix, and can be exactly computed in polynomial time. We thus denote by $\text{Lip}(\text{FC}_{1})$ and $\text{Lip}(\text{FC}_{2})$ the Lipschitz constants of the two fully-connected layers, respectively. Therefore, we can compute an upper bound of the Lipschitz constant of some models using Theorem 3.3 and Lemma 3.4. To train the models, we add a final layer to the models which transforms the vector representations of the multisets into class probabilities.
We use the same experimental protocol as in subsection 4.1 above (i. e., we randomly choose $100$ test samples). We only provide results for ModelNet40 in Figure 2, while the results for Polarity can be found in Appendix D.2. Since all point clouds contained in the ModelNet40 dataset have equal cardinalities, the results of both Theorem 3.3 and Lemma 3.4 hold, and thus, once again, we can derive upper bounds on the Lipschitz constants for $7$ out of the $9$ combinations of distance functions for multisets and aggregation functions. We observe that the dash lines (Lipschitz upper bounds from Theorem 3.3 and Lemma 3.4) indeed upper bound the Euclidean distance of the outputs of the aggregation functions. We can also see that the bounds that are associated with the mean function are tight, while those associated with the sum and max functions are relatively loose and very loose, respectively. We also observe that the distances of the representations produced by the mean and sum functions are very correlated with the distances produced by EMD and matching distance. On the other hand, the max function gives rise to representations whose distances are less correlated with the distances produced by the distance functions for multisets.
@@ 4.3 Stability under Perturbations of Input Multisets
We now empirically study the stability of the two Lipschitz continuous models ($\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$) under perturbations of the input multisets. Our objective is to apply small perturbations to test samples such that the models misclassify the perturbed samples.
@@
We consider two different perturbations, Pert. #1 and Pert. #2. Both perturbations are applied to test samples once each model has been trained. We then examine whether the perturbation leads to a decrease in the accuracy achieved on the test set. Pert. #1 is the perturbation described in Proposition 3.5 and is applied to the multisets of ModelNet40. Specifically, we add to each test sample a single element. We choose to add the element that has the highest norm across the elements of all samples. Pert. #2 is applied to the multisets of Polarity. It adds random noise to each element of each multiset of the test set. Specifically, a random vector is sampled from $\mathcal{U}(0,0.2)^{d}$ and is added to each element of each test sample. The results are provided in Table 2. $\textsc{NN}_{\textsc{max}}$ appears to be insensitive to Pert. #1, while $\textsc{NN}_{\textsc{mean}}$ is insensitive to Pert. #2. The results indicate that $\textsc{NN}_{\textsc{mean}}$ is more robust than $\textsc{NN}_{\textsc{max}}$ to a larger perturbation that is associated with a single or a few elements of the multiset. On the other hand, $\textsc{NN}_{\textsc{max}}$ is more robust to smaller perturbations applied to all elements of the multiset.
Figure 3: Size Generalization of $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models. For illustration purposes, the Wasserstein distances $\mathcal{W}_{1}$ are normalized to make the maximal distance equal to the greatest performance drops. The models of the left plots are trained on the first bucket, while those of the right plots on the last bucket.
@@ 4.4 Generalization under Distribution Shifts
Finally, we investigate whether neural networks for multisets can generalize to multisets of different cardinalities. We randomly sample $2,000$ documents from the Polarity dataset, and we represent them as multisets of word vectors. We then sort the multisets of word vectors based on their cardinality, and construct $10$ bins, each containing $200$ multisets. The $i$-th bin contains multisets $X_{(200i)+1},X_{(200i)+2},\ldots,X_{(200i)+200}$ from the sorted list of multisets. We then train $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ (which are Lipschitz continuous) on the first and the last bin and once the models are trained, we compute their accuracy on all $10$ bins. We also compute the Wasserstein distance with $p=1$ between domain distributions (i. e., between the first bin and the rest of the bins, and also between the last bin and the rest of the bins). We then aim to validate Theorem 3.6 which states that the error on different domains can be bounded by the Wasserstein distance between the data distributions. We thus compute the correlation between tha accuracy drop and the Wasserstein distance between the two distributions. The results for $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ are illustrated in Figure 3. The results are averaged over $10$ runs.
We observe that the Wasserstein distance between the data distributions using EMD (for $\textsc{NN}_{\textsc{mean}}$) and Hausdorff distance (for $\textsc{NN}_{\textsc{max}}$) as ground metrics highly correlates with the accuracy drop both in the case where the $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models are trained on small multisets and tested on larger multisets ($r=0.92$ and $r=0.90$, respectively) and also in the case where the models are trained on large multisets and tested on smaller multisets ($r=0.94$ and $r=0.90$, respectively). The correlation is slightly weaker in the case of the $\textsc{NN}_{\textsc{max}}$ model. Our results suggest that the drop in accuracy is indeed related to the Wasserstein distance between the data distributions, and that it can provide insights into the generalization performance of the models under distribution shifts.
5 Conclusion
In this paper, we focused on data modeled as multisets of vectors and found that for multisets that differ from each other in cardinality, each aggregation function is Lipschitz continuous with respect to only one of the three distance functions. We have also computed the Lipschitz constants of those functions. We then derived upper bounds on the Lipschitz constants of neural networks that process multisets of vectors, and studied their stability to perturbations and generalization under distribution shifts. Our theoretical results were confirmed by numerical experiments.
Limitations.
The Lipschitz constants of aggregation functions and upper bounds of Lipschitz constants of neural networks might not be tight in practice. Specifically, our experiments demonstrate that the Lipschitz constant of the max function is not tight in the case of the ModelNet40 and Polarity datasets. The actual data samples that can result into a very tight bound might be far from realistic. Likewise, the derived upper bounds of the Lipschitz constants of $\textsc{NN}_{\textsc{mean}}$, $\textsc{NN}_{\textsc{sum}}$ and $\textsc{NN}_{\textsc{max}}$ do not seem to be very tight in most cases.
Appendix A Proofs
@@
$d_{M}(X,Y)=\min_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||+\sum_{i=n+1}^{m}||\mathbf{v}_{\pi(i)}||\bigg{]}\geq\sum_{i=n+1}^{m}||\mathbf{v}_{\pi(i)}||>0$
@@
since $||\mathbf{v}||>0$ for all $\mathbf{v}\in\mathbb{R}^{d}\setminus\{\mathbf{0}\}$. If $|X|=|Y|=m$, since the two multisets are different from each other, there exists at least one vector $\mathbf{u}_{i}$ with $i\in[m]$ such that $\mathbf{u}_{i}\in Y$, but $\mathbf{u}_{i}\not\in X$. Let $\pi^{*}\in\mathfrak{S}_{m}$ denote the permutation associated with $d_{M}(X,Y)$. Then, we have that:
$d_{M}(X,Y)=||\mathbf{v}_{\pi^{*}(1)}-\mathbf{u}_{1}||+\ldots+||\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}||+\ldots+||\mathbf{v}_{\pi^{*}(m)}-\mathbf{u}_{m}||\geq||\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}||>0$
@@
Thus, we only need to prove that the triangle inequality holds. Let $S\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$. The three multisets can have different cardinalities. Let $|X|=m$, $|Y|=n$ and $|S|=k$. Then, $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{n}\}$ and $Z=\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{k}\}$. There are $6$ different cases. But it suffices to show that the triangle inequality holds when $|X|\geq|Y|\geq|Z|$, when $|Z|\geq|Y|\geq|X|$ and when $|X|\geq|Z|\geq|Y|$. Proofs for the rest of the cases are similar.
@@
Case 1: Suppose $|X|\geq|Y|\geq|Z|$. Let $S_{1}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(X,Z)$
$\pi_{1}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{k}||\mathbf{v}_{\pi(i)}-\mathbf{z}_{i}||+\sum_{i=k+1}^{m}||\mathbf{v}_{\pi(i)}||\bigg{]}$
Likewise, let $\pi_{2}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(Z,Y)$
$\pi_{2}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{n}}\bigg{[}\sum_{i=1}^{k}||\mathbf{u}_{\pi(i)}-\mathbf{z}_{i}||+\sum_{i=k+1}^{n}||\mathbf{u}_{\pi(i)}||\bigg{]}$
Then, we have
$d_{M}(X,Z)+d_{M}(Z,Y)$ $=\sum_{i=1}^{k}||\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{z}_{i}||+\sum_{i=k+1}^{m}||\mathbf{v}_{\pi_{1}^{*}(i)}||+\sum_{i=1}^{k}||\mathbf{z}_{i}-\mathbf{u}_{\pi_{2}^{*}(i)}||+\sum_{i=k+1}^{n}||\mathbf{u}_{\pi_{2}^{*}(i)}||$
$=\sum_{i=1}^{k}\Big{[}||\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{z}_{i}||+||\mathbf{z}_{i}-\mathbf{u}_{\pi_{2}^{*}(i)}||\Big{]}+\sum_{i=k+1}^{n}\Big{[}||\mathbf{v}_{\pi_{1}^{*}(i)}||+||-\mathbf{u}_{\pi_{2}^{*}(i)}||\Big{]}+\sum_{i=n+1}^{m}||\mathbf{v}_{\pi_{1}^{*}(i)}||$
$\geq\sum_{i=1}^{k}||\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{u}_{\pi_{2}^{*}(i)}||+\sum_{i=k+1}^{n}||\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{u}_{\pi_{2}^{*}(i)}||+\sum_{i=n+1}^{m}||\mathbf{v}_{\pi_{1}^{*}(i)}||$
$=\sum_{i=1}^{n}||\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{u}_{\pi_{2}^{*}(i)}||+\sum_{i=n+1}^{m}||\mathbf{v}_{\pi_{1}^{*}(i)}||$
$\geq\min_{\pi\in\mathfrak{S}_{m}}\bigg{[}\sum_{i=1}^{n}||\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}||+\sum_{i=n+1}^{m}||\mathbf{v}_{\pi(i)}||\bigg{]}$

### v3

@@ 4.2 Upper Bounds of Lipschitz Constants of Neural Networks for Sets
In the second set of experiments, we empirically validate the results of Theorem 3.4 and Lemma 3.5 on the ModelNet40 and Polarity datasets. We build neural networks that consist of three layers: (i) a fully-connected layer; (ii) an aggregation function; and (iii) a second fully-connected layer. Therefore, those models first transform the elements of the input multisets using an affine function, then aggregate the representations of the elements of each multiset and finally they transform the aggregated representations using a another affine function. Note that the Lipschitz constant of an affine function is equal to the largest singular value of the associated weight matrix, and can be exactly computed in polynomial time. We thus denote by $\text{Lip}(\text{FC}_{1})$ and $\text{Lip}(\text{FC}_{2})$ the Lipschitz constants of the two fully-connected layers, respectively. Note also that the Lipschitz constant of most activation functions (e. g., ReLU, LeakyReLU, Tanh) is equal to $1$. Therefore, we can compute an upper bound of the Lipschitz constant of some models using Theorem 3.4 and Lemma 3.5. To train the models, we add a final layer to them which transforms the vector representations of the multisets into class probabilities. We use the same experimental protocol as in subsection 4.1 above (i. e., we randomly choose $100$ test samples). We only provide results for ModelNet40 in Figure 2, while the results for Polarity can be found in Appendix E.2. Since all point clouds contained in the ModelNet40 dataset have equal cardinalities, the conclusions of both Theorem 3.4 and Lemma 3.5 apply to this setting, and thus, once again, we can derive upper bounds on the Lipschitz constants for $7$ out of the $9$ combinations of distance functions for multisets and aggregation functions. We observe that the dash lines (Lipschitz upper bounds from Theorem 3.4 and Lemma 3.5) indeed upper bound the Euclidean distance of the outputs of the aggregation functions. We can also see that the bounds that are associated with the mean function are tight, while those associated with the sum and max functions are relatively loose and very loose, respectively. We also observe that the distances of the representations produced by the mean and sum functions are very correlated with the distances produced by EMD and matching distance. On the other hand, the max function gives rise to representations whose distances are less correlated with the distances produced by the distance functions for multisets.
@@ 4.3 Stability under Perturbations of Input Multisets
We now empirically study the stability of the two Lipschitz continuous models ($\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$) under perturbations of the input multisets. Our objective is to apply small perturbations to test samples such that the models misclassify the perturbed samples. We consider two different perturbations, Pert. #1 and Pert. #2. Both perturbations are applied to test samples once each model has been trained. We then examine whether the perturbation leads to a decrease in the accuracy achieved on the test set. Pert. #1 is the perturbation described in Proposition 3.6 and is applied to the multisets of ModelNet40. Specifically, we add to each test sample a single element. We choose to add the element that has the highest norm across the elements of all samples. Pert. #2 is applied to the multisets of Polarity. It adds random noise to each element of each multiset of the test set. Specifically, a random vector is sampled from $\mathcal{U}(0,0.2)^{d}$ and is added to each element of each test sample. We selected this distribution because its mean and standard deviation closely match the empirical mean and standard deviation of the dataset’s word vectors. The results are provided in Table 2. $\textsc{NN}_{\textsc{mean}}$ appears to be insensitive to Pert. #1, while $\textsc{NN}_{\textsc{max}}$ is insensitive to Pert. #2. The results indicate that $\textsc{NN}_{\textsc{mean}}$ is more robust than $\textsc{NN}_{\textsc{max}}$ to a larger perturbation that is associated with a single or a few elements of the multiset. On the other hand, $\textsc{NN}_{\textsc{max}}$ is more robust to smaller perturbations applied to all elements of the multiset.
Figure 3: Size Generalization of $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models. For illustration purposes, the Wasserstein distances $\mathcal{W}_{1}$ are normalized to make the maximal distance equal to the greatest performance drops. The models in the left plots are trained on the first bucket, while those in the right plots are trained on the last bucket.
@@ 4.4 Generalization under Distribution Shifts
Finally, we investigate whether neural networks for multisets can generalize to multisets of different cardinalities. We randomly sample $2,000$ documents from the Polarity dataset, and we represent them as multisets of word vectors. We then sort the multisets of word vectors based on their cardinality, and construct $10$ bins, each containing $200$ multisets. The $i$-th bin contains multisets $X_{(200\cdot i)+1},X_{(200\cdot i)+2},\ldots,X_{(200\cdot i)+200}$ from the sorted list of multisets. We then train $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ (which are Lipschitz continuous) on the first and the last bin and once the models are trained, we compute their accuracy on all $10$ bins. We also compute the Wasserstein distance with $p=1$ between domain distributions (i. e., between the first bin and the rest of the bins, and also between the last bin and the rest of the bins). We then aim to validate Theorem 3.7 which states that the error on different domains can be bounded by the Wasserstein distance between the data distributions. We thus compute the correlation between the accuracy drop and the Wasserstein distance between the two distributions. The results for $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ are illustrated in Figure 3. The results are averaged over $10$ runs. We observe that the Wasserstein distance between the data distributions using EMD (for $\textsc{NN}_{\textsc{mean}}$) and Hausdorff distance (for $\textsc{NN}_{\textsc{max}}$) as ground metrics highly correlates with the accuracy drop both in the case where the $\textsc{NN}_{\textsc{mean}}$ and $\textsc{NN}_{\textsc{max}}$ models are trained on small multisets and tested on larger multisets ($r=0.92$ and $r=0.90$, respectively) and also in the case where the models are trained on large multisets and tested on smaller multisets ($r=0.94$ and $r=0.90$, respectively). The correlation is slightly weaker in the case of the $\textsc{NN}_{\textsc{max}}$ model. Our results suggest that the drop in accuracy is indeed related to the Wasserstein distance between the data distributions, and that it can provide insights into the generalization performance of the models.
5 Conclusion and Discussion
In this paper, we studied the Lipschitz continuity of multiset aggregation functions with respect to three distance functions. We also explored the Lipschitz constants of neural networks that process multisets of vectors. As a general guideline, one should choose the aggregation function that is Lipschitz continuous with respect to the distance function that best captures the distances between the multisets in the considered dataset or problem. For example, in problems where the shape of the input object matters (e. g., shapes extracted from medical images or 3D scans), Hausdorff distance is preferable to EMD and the matching distance since we would like to detect whether any part of one shape is far away from the other shape, even if the rest of the shapes are well-aligned. However, in some cases, not a single distance function is suitable for a single problem. For instance, consider the problem of text categorization, where documents are represented as multisets of word vectors. If two documents are considered similar when they contain similar terms, regardless of their length, the EMD is likely to best capture the distance between them. On the other hand, if similarity is determined by the presence of just one or a few extreme shared words, the Hausdorff distance is more appropriate. This illustrates that selecting an aggregation function typically requires some domain knowledge. In the absence of such knowledge, choosing an aggregation function can be challenging, except in special cases, such as when multisets have the same cardinality where our results indicate that the max function is Lipschitz continuous with respect to all distance functions.
Acknowledgements
We thank the anonymous reviewers for their helpful comments and suggestions. This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Appendix A Related Work
In recent years, there has been an increasing interest in applying machine learning algorithms to set-structured data. Since sets and multisets are inherently unordered, these models need to be invariant to permutations of the elements in the input set. It has been shown that, otherwise, the ordering of inputs strongly affects performance (Vinyals et al., 2016).
The seminal work of Zaheer et al. (2017) introduced DeepSets, a model that uses the sum aggregator to produce permutation-invariant set representations. They showed that when the elements of the input sets come from a countable set $\mathcal{X}$, then for an appropriate $f\colon\mathcal{X}\rightarrow\mathbb{R}$, the function defined as $g(\{x_{1},\ldots,x_{n}\})=\sum_{i=1}^{n}f(x_{i})$ maps the input sets injectively to $\mathbb{R}$. This result was later extended to multisets (Xu et al., 2019). In the countable case, an embedding dimension of $1$ already suffices for injectivity. When $\mathcal{X}=\mathbb{R}$, multiset cardinalities are bounded by $m$ and $f$ is continuous, an embedding dimension of at least $m$ is both necessary and sufficient for injectivity (Wagstaff et al., 2019; 2022). For $\mathcal{X}=\mathbb{R}^{d}$, an embedding dimension of at least $md$ is necessary (Joshi et al., 2023). Most of these results rely on polynomial constructions to build injective multiset functions, after which the universal approximation theorem is invoked to argue that MLPs can approximate such polynomials. Amir et al. (2023) investigate whether MLP-based multiset functions are actually injective. They show that injectivity depends on the activation function. Specifically, analytic non-polynomial activation functions yield injective models, while networks with piecewise linear activation functions are injective only when $\mathcal{X}$ is finite or corresponds to certain irregular, countably infinite sets.
Besides DeepSets, several other architectures and aggregation functions have been proposed recently. PointNet is another important architecture, primarily designed for point cloud data (Qi et al., 2017a). It consists of the same components as DeepSets, but instead of a sum aggregation function, it employs a max aggregator. To allow PointNet to capture local structures at different scales, Qi et al. (2017b) proposed PointNet++, a hierarchical model which applies PointNet recursively to nested partitions of the input set. Janossy pooling applies a neural network to all permutations of the input data and averages their outputs (Murphy et al., 2019). Since computing all permutations is generally intractable, the authors also propose some practical approximations. Set Transformer is a variant of the Transformer architecture designed for sets (Lee et al., 2019). Due to its attention mechanism, the model can capture interactions between elements in the input set. RepSet is another model designed for set-structured data which generates set representations by comparing the input set against learnable latent sets using a network flow algorithm (Skianis et al., 2020). FSPool sorts each feature across the elements of the set, and then computes a weighted sum of the elements where different weights can be learned for each feature dimension (Zhang et al., 2020). Pellegrini et al. (2021) propose a a learnable aggregation function which can approximate common aggregation functions (e. g., mean, sum, max), but also more complex functions. Kimura et al. (2024) introduce the Hölder’s Power DeepSets, a model tha generalized DeepSets by employing function known as power‑mean (a.k.a. Hölder mean), controlled by an exponent $p$.
Recently, a line of works has studied the Lipschitz continuity of aggregation functions and has developed embeddings that are bi-Lipschitz. Amir et al. (2023) showed that although DeepSets models that use analytic non-polynomial activation functions are injective, they are not bi-Lipschitz with respect to the 2-Wasserstein distance. Davidson and Dym (2024) investigated the Hölder continuity of neural networks for sets, a relaxation of Lipschitz continuity. They relied on a probabilistic framework of Hölder stability in expectation and showed that DeepSets with ReLU activation functions have an expected lower-Hölder exponent of $3/2$, whereas smooth activation functions yield a much worse expected lower-Hölder exponent. Balan et al. (2025) presented an embedding scheme based on sorting random projections of the multiset elements. The embedding is shown to be injective and bi-Lipschitz. The Fourier Sliced–Wasserstein (FSW) embedding is another theoretically grounded method for learning representations of sets (Amir and Dym, 2025). It computes random projections of the input data, and for each projection it samples the cosine transform of the corresponding quantile function. From a theoretical standpoint, the FSW embedding has a significant advantage over most previous methods, as it is proven to be both injective and bi-Lipschitz.
Neural network models that can handle set-structured data have been applied across diverse domains, including biology (Clarke et al., 2024), chemistry (Boulougouri et al., 2024) and materials science (Zhang et al., 2022b). In some applications, domain knowledge is incorporated into those models. For instance, Lim et al. (2023) introduce neural architectures specifically designed for eigenvector-based inputs, which can be viewed as variants of DeepSets, while explicitly accounting for the symmetries inherent in eigenvectors. For a recent overview of neural network models for set-structured data, we refer the reader to the survey by Xie and Tong (2025).
Appendix B Proofs
@@
$d_{M}(X,Y)=\min_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|+\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi(i)}\|\bigg]\geq\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi(i)}\|>0$
@@
since $\|\mathbf{v}\|>0$ for all $\mathbf{v}\in\mathbb{R}^{d}\setminus\{\mathbf{0}\}$. If $|X|=|Y|=m$, since the two multisets are different from each other, there exists at least one vector $\mathbf{u}_{i}$ with $i\in[m]$ such that $\mathbf{u}_{i}\in Y$, but $\mathbf{u}_{i}\not\in X$. Let $\pi^{*}\in\mathfrak{S}_{m}$ denote a permutation associated with $d_{M}(X,Y)$. Then, we have that:
$d_{M}(X,Y)=\|\mathbf{v}_{\pi^{*}(1)}-\mathbf{u}_{1}\|+\ldots+\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|+\ldots+\|\mathbf{v}_{\pi^{*}(m)}-\mathbf{u}_{m}\|\geq\|\mathbf{v}_{\pi^{*}(i)}-\mathbf{u}_{i}\|>0$
@@
Thus, we only need to prove that the triangle inequality holds. Let $Z\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$. The three multisets can have different cardinalities. Let $|X|=m$, $|Y|=n$ and $|Z|=k$. Then, $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}\!\!\}$, $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{n}\}\!\!\}$ and $Z=\{\!\!\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{k}\}\!\!\}$. There are $6$ different cases. But it suffices to show that the triangle inequality holds when $|X|\geq|Y|\geq|Z|$, when $|Z|\geq|Y|\geq|X|$ and when $|X|\geq|Z|\geq|Y|$. Proofs for the rest of the cases are similar.
@@
Case 1: Suppose $|X|\geq|Y|\geq|Z|$. Let $\pi_{1}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(X,Z)$:
$\pi_{1}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{k}\|\mathbf{v}_{\pi(i)}-\mathbf{z}_{i}\|+\sum_{i=k+1}^{m}\|\mathbf{v}_{\pi(i)}\|\bigg]$
Likewise, let $\pi_{2}^{*}$ denote the matching produced by the solution of the matching distance function $d_{M}(Z,Y)$:
$\pi_{2}^{*}=\operatorname*{arg\,min}_{\pi\in\mathfrak{S}_{n}}\bigg[\sum_{i=1}^{k}\|\mathbf{u}_{\pi(i)}-\mathbf{z}_{i}\|+\sum_{i=k+1}^{n}\|\mathbf{u}_{\pi(i)}\|\bigg]$
Then, we have:
$d_{M}(X,Z)+d_{M}(Z,Y)$ $=\sum_{i=1}^{k}\|\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{z}_{i}\|+\sum_{i=k+1}^{m}\|\mathbf{v}_{\pi_{1}^{*}(i)}\|+\sum_{i=1}^{k}\|\mathbf{z}_{i}-\mathbf{u}_{\pi_{2}^{*}(i)}\|+\sum_{i=k+1}^{n}\|\mathbf{u}_{\pi_{2}^{*}(i)}\|$
$=\sum_{i=1}^{k}\Big[\|\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{z}_{i}\|+\|\mathbf{z}_{i}-\mathbf{u}_{\pi_{2}^{*}(i)}\|\Big]+\sum_{i=k+1}^{n}\Big[\|\mathbf{v}_{\pi_{1}^{*}(i)}\|+\|-\mathbf{u}_{\pi_{2}^{*}(i)}\|\Big]+\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi_{1}^{*}(i)}\|$
$\geq\sum_{i=1}^{k}\|\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{u}_{\pi_{2}^{*}(i)}\|+\sum_{i=k+1}^{n}\|\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{u}_{\pi_{2}^{*}(i)}\|+\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi_{1}^{*}(i)}\|$
$=\sum_{i=1}^{n}\|\mathbf{v}_{\pi_{1}^{*}(i)}-\mathbf{u}_{\pi_{2}^{*}(i)}\|+\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi_{1}^{*}(i)}\|$
$\geq\min_{\pi\in\mathfrak{S}_{m}}\bigg[\sum_{i=1}^{n}\|\mathbf{v}_{\pi(i)}-\mathbf{u}_{i}\|+\sum_{i=n+1}^{m}\|\mathbf{v}_{\pi(i)}\|\bigg]$